Find the measure of an angle that is 30 degrees less than its supplement
Answer by checkley77(12569)

x+(x-30)=180
2x=180-30
2x=150
x=150/2
x=75 degrees ans. for the smaller angle.
180-75=105 degrees for the larger angle.